OpenStudy (anonymous):

Algebra 2/Trig Simplify 4^1.5 * 2^-2

OpenStudy (anonymous):

\[4=2^2\] so \[4^{1.5}=2^3\]

OpenStudy (anonymous):

now add the exponents to finish it